# Efficient Tool Use with Chain-of-Abstraction Reasoning

### 30 Jan 2024 v1

@@ Appendix A Implementation Details
For mathematical reasoning evaluation, we extract the last number appeared in each model’s answer, and check whether the number exactly match the gold reference. The accuracy is reported as the rate of such exact match across all QAs in a test set. For Wiki QA evaluation, similar to mathematical reasoning, we extract the final answer of each model and calculate its exact match rate to the gold reference. Specifically, the final answer is supposed to be the words after “Action: finish[” for FireAct baseline, and words after “The answer is ” for other models. Our 8-shot in-domain examples used for the CoT-FSP baseline are shown in Table 13 and 14, which enables the model to provide answer with our required format for evaluation, i.e., stating its final answer after “The answer is ”. Our human evaluation on GSM8K is conducted by 5 internal domain experts from our research group. For each math question, we provide the experts with the gold answer as reference, and ask them to evaluate each model answer in anonymous manner, i.e., experts do not know which model each answer comes from. Two yes-or-no questions are asked for evaluating each model answer, including: a) whether the answer has any arithmetic error, and b) whether the answer has any reasoning error, and binary choices from the experts are collected to calculate the error rates of each model’s generation. We present our detailed instructions for human evaluation in Figure 5.
@@
We fine-tune our models with batch size $8$ and learning rate $2e^{-5}$ and $1e^{-5}$ for 7B and 70B model sizes, respectively, using cosine learning rate scheduler with warm-up step $10$. We use AdamW (Loshchilov and Hutter, 2018) optimizer for all our fine-tuning experiments, with $\beta_{1}$, $\beta_{2}$ and $\epsilon$ set to $0.9$, $0.95$ and $1e^{-8}$, respectively. Training weight decay is set to $0.1$. For mathematical reasoning, we use a total of $400$ training steps, and get the best model checkpoints (with highest validation scores) at step $240$ and $200$ for 7B and 70B model sizes. For Wiki QA domain, we adjust the total training steps to $500$, and get the best checkpoints at step $450$ and $300$ for 7B and 70B models. Therefore, only $\sim$2K and $\sim$3K QAs are required in practice for fine-tuning our models in math and Wiki QA domains. The training of our 7B and 70B models is based on 8 and 64 NVIDIA A100-SXM4 (80GB) GPUs, respectively.
@@ Appendix B Full Experimental Results
Table 9 and 10 show the full results of our experiments on math and Wiki QA domains. Our method of CoA achieves consistent improvements over baselines across various LLaMa model versions (LLaMa, LLaMa-2 and LLaMa-2-Chat), model sizes (7B and 70B), and domain benchmarks. This shows great potential of our method being generalized to new model backbones and reasoning tasks.

### 26 Feb 2024 v2

@@ Appendix A Implementation Details
For mathematical reasoning evaluation, we extract the last number appeared in each model’s answer, and check whether the number exactly match the gold reference. The accuracy is reported as the rate of such exact match across all QAs in a test set. For Wiki QA evaluation, similar to mathematical reasoning, we extract the final answer of each model and calculate its exact match rate to the gold reference. Specifically, the final answer is supposed to be the words after “Action: finish[” for FireAct baseline, and words after “The answer is ” for other models. Our 8-shot in-domain examples used for the CoT-FSP baseline are shown in Table 13 and 14, which enables the model to provide answer with our required format for evaluation, i.e., stating its final answer after “The answer is ”. Our human evaluation on GSM8K is conducted by 5 internal domain experts from our research group. For each math question, we provide the experts with the gold answer as reference, and ask them to evaluate each model answer in anonymous manner, i.e., experts do not know which model each answer comes from. Two yes-or-no questions are asked for evaluating each model answer, including: a) whether the answer has any arithmetic error, and b) whether the answer has any reasoning error, and binary choices from the experts are collected to calculate the error rates of each model’s generation. We present our detailed instructions for human evaluation in Figure 5. Our data collection protocol is approved by our organization in terms of ethics.
@@
We fine-tune our models with batch size $8$ and learning rate $2e^{-5}$ and $1e^{-5}$ for 7B and 70B model sizes, respectively, using cosine learning rate scheduler with warm-up step $10$. We use AdamW (Loshchilov and Hutter, 2018) optimizer for all our fine-tuning experiments, with $\beta_{1}$, $\beta_{2}$ and $\epsilon$ set to $0.9$, $0.95$ and $1e^{-8}$, respectively. Training weight decay is set to $0.1$. For mathematical reasoning, we use a total of $400$ training steps, and get the best model checkpoints (with highest validation scores) at step $240$ and $200$ for 7B and 70B model sizes. For Wiki QA domain, we adjust the total training steps to $500$, and get the best checkpoints at step $450$ and $300$ for 7B and 70B models. Therefore, only $\sim$2K and $\sim$3K QAs are required in practice for fine-tuning our models in math and Wiki QA domains. The training of our 7B and 70B models is based on 8 and 64 NVIDIA A100-SXM4 (80GB) GPUs, with training time about 2 and 5 hours per model, respectively.
@@ Appendix B Full Experimental Results
Table 7 and 8 show the full results of our experiments on math and Wiki QA domains. Our method of CoA achieves consistent improvements over baselines across various LLaMa model versions (LLaMa, LLaMa-2 and LLaMa-2-Chat), model sizes (7B and 70B), and domain benchmarks. This shows great potential of our method being generalized to new model backbones and reasoning tasks. We also present results on GSM8K subsets according to varying numbers of gold reasoning steps in Table 9, where we confirm that CoA has more robust long chain-of-thought reasoning accuracy.
@@
Method Gold Reasoning Step
$\leq 2$ $3$ $4$ $5$ $>5$
CoT-FSP 42.9 26.3 18.0 10.9 3.6
CoT-FT 55.5 42.6 25.8 19.0 10.8
CoA 55.8 44.4 32.5 25.3 15.1
+0.3 +1.8 +6.7 +6.3 +4.3
Method Accuracy
CoT-FSP 27.90
CoT-FT 39.12
Toolformer 24.56
Toolformer - Math 35.25
CoA 40.79
@@ Appendix C Fine-Tuning Data Re-writing Details
Appendix D Claim of Usage
Our use of existing scientific artifacts cited in this paper is consistent with their intended use. Our developed code, data and models are intended to be used for only research purposes, any usage of our scientific artifacts that is outside of research contexts should not be allowed.